y3-xy2-14x2y=24x3=0......solve this equation for getting three separate straight lines....ans not at all matching...

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear shivanshi

y3-xy2-14x2y+24x3=0

y3-2xy2 + xy2-2x2y-12x2y+24x3=0

y2(y-2x) +xy(y-2x) -12x2(y-2x)=0

(y-2x)(y2 +xy -12x2)=0

(y-2x)(y-3x)(y+4x)=0

three lines are

y=2x

y=3x

y=-4x

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin